# Markov switching zero-inflated space-time multinomial models for comparing multiple infectious diseases

### v1

@@ 4.1 Model specification and fitting
We modeled the Rio data with a bi-weekly time step as this roughly corresponds to the serial interval of the three diseases (Majumder et al., 2016; Riou et al., 2017). As mentioned in Section 2.1, we took dengue as the baseline disease (k=1) since it is well-established in the city and cases are reported consistently throughout the study period. We then considered $k=2$ to represent Zika and $k=3$ to represent chikungunya.
@@
For $\bm{x}_{kit}$ in (3), covariates that may affect the favorability of Zika or chikungunya transmission relative to dengue transmission, we first took all spatial covariates considered in Schmidt et al. (2022): the percentage of neighborhood $i$ covered in green area, $\text{verde}_{i}$; the social development index of neighborhood $i$, $\text{SDI}_{i}$; and the population density of neighborhood $i$, $\text{popdens}_{i}$. We additionally considered the percentage of neighborhood $i$ that was occupied by favelas (slums), $\text{favela}_{i}$. These spatial covariates were chosen as the transmission of arboviral diseases can be significantly influenced by socioeconomic factors such as water supply and urbanicity (Schmidt et al., 2011), see Schmidt et al. (2022) for more details. As for space-time covariates, we first took the average weekly maximum temperature in neighborhood $i$ during bi-week $t$, $\text{temp}_{it}$. Temperature is an important factor to consider since if the transmission of Zika or chikungunya is more sensitive (less sensitive) to temperature compared to dengue then their range of spread will be smaller (bigger) (Tesla et al., 2018; Mercier et al., 2022). Note, if there are many more cases of Zika or chikungunya in neighboring areas, compared to dengue, then we would expect the share of those diseases to grow all else equal, due to between-area mixing (Stoddard et al., 2013). Therefore, we considered the previous prevalence of disease $k$ and disease $1$ (dengue) across neighboring areas, $\log\mathopen{}\mathclose{{}\left(\sum_{j\in NE(i)}y_{kj(t-1)}/\sum_{j\in NE(i)}\text{pop}_{j}+1}\right)$ and $\log\mathopen{}\mathclose{{}\left(\sum_{j\in NE(i)}y_{1j(t-1)}/\sum_{j\in NE(i)}\text{pop}_{j}+1}\right)$, in $\bm{x}_{kit}$, where $NE(i)$ is the set of neighboring areas of neighborhood $i$ and $\text{pop}_{j}$ is the population of neighborhood $j$. We considered two areas neighbors if they shared a border. To simplify the model, we assumed the effect of neighboring dengue prevalence was the same on both $\log(\pi_{2it}/\pi_{1it})$ and $\log(\pi_{3it}/\pi_{1it})$, like with the within-area autoregression. Finally, it was speculated in Freitas et al. (2019) that Zika circulation could have been inhibiting chikungunya transmission. To account for these kinds of disease interactions, we considered $\log(y_{3i(t-1)}+1)$ in $\bm{x}_{2it}$ and $\log(y_{2i(t-1)}+1)$ in $\bm{x}_{3it}$. That is, chikungunya cases were allowed to affect the favorability of Zika transmission relative to dengue transmission and Zika cases were allowed to affect the favorability of chikungunya transmission relative to dengue transmission.
@@
As for covariates potentially associated with the presence of Zika or chikungunya, $\bm{z}_{kit}$ in (7), we considered mostly the same factors included in $\bm{x}_{kit}$. An exception is that we did not include the cases of the other disease as the Markov chain defined in (7) already accounts for disease interactions through $\rho_{jk}^{DI}$. Also, since the presence is not defined relatively, we did not include neighboring dengue prevalence.
@@
We also need to specify the initial state distributions, see Section 3. As there were no chikungunya cases reported at all in the first year, we assumed chikungunya had only a 5 percent chance of being present at the start of 2015. For Zika, there are a few, but not many, cases reported through most of 2015 and, therefore, we assumed a 10 percent chance of initial presence. We fitted the MS-ZIARMN model specified above to the Rio data using our proposed Gibbs sampler from Section 3. We ran the Gibbs sampler for 250,000 iterations on three chains with an initial burn-in of 50,$0$ iterations. All sampling was started from random values in the parameter space to avoid convergence to local modes. Convergence was checked using the Gelman-Rubin statistic ($<$1.05), the minimum effective sample size ($>$1000) and by visually examining the traceplots (Plummer et al., 2006).
@@
For comparison purposes, we also considered a model without zero inflation to ensure there is justification for treating the many zeroes in the data, as hypothesized. That is, the ARMN model described in Equations (1)-(4) with the same covariates $\bm{x}_{2it}$ and $\bm{x}_{3it}$ as the MS-ZIARMN model. We also considered a zero-inflated model without the coupled Markov chains, that is, with $\rho_{k}^{AR}=0$ and $\rho_{jk}^{DI}=0$ for all $k$ and $j>1,\,j\neq k$ in Equation (7), which we will call the ZIARMN model. Without the coupled Markov chains, our model is a finite mixture and not a Markov mixture model, and the inference is greatly simplified, e.g., there is no need to use the FFBS algorithm. Finally, we compared with the zero-inflated multinomial model of Zeng et al. (2022) as our approach is built on their framework. Zeng et al. (2022) assumed that the probability a category was absent from an observation only depended on the category, so we let $\text{logit}(p_{kit})=\eta_{0k}$ for $k=2,3$ in Equation (7). We kept the rest of the model the same as the MS-ZIARMN model. Note, Zeng et al. (2022) did not consider any covariates or space-time structure in the multinomial probabilities. However, this is inappropriate for our example and did not fit the data well (results not shown). Also, Zeng et al. (2022) did not assume the baseline category was always present, however, this appears to be necessary when modeling a small number of categories, see Section 2.1.
@@
Table 1 shows the widely applicable information criterion (WAIC) (Gelman et al., 2014) of the 4 considered models. We calculated the WAIC by marginalizing out the state indicators $\bm{S}^{*}$ as recommended by Auger-Méthé et al. (2021), see SM Section 4 for more details. Note, the model with the smallest WAIC is considered to have the best fit and as a rule of thumb a difference of 5 or more in the WAIC is considered significant. Table 1 shows it is not only important to account for zero inflation, but also covariates and correlations across space, time and disease in the zero-inflated process (i.e., the process of disease presence). Indeed, the model of Zeng et al. (2022) had a worse fit compared to the ARMN model (without zero inflation), despite the other zero-inflated models having a much better fit.
@@ 4.2 Results
Table 2 shows the estimated coefficients from the multinomial part of the fitted MS-ZIARMN model, that is the intercepts and covariate effects from Equation (3). From the intercept row, at average values of the covariates, Zika transmission was favored over dengue and chikungunya transmission. That is, under average conditions the share of Zika cases tended to grow over time in an area (at least from an equal share, see the nuances in Section 2). It is interesting to compare these estimated intercepts with those from the ARMN and Zeng (2022) models in SM Table 1. Unlike the MS-ZIARMN model, the Zeng (2022) and ARMN models both estimated that on average Zika and chikungunya transmission was much less intense in Rio compared to dengue transmission. This does not correspond to our knowledge of the epidemiology of these diseases and is likely due to a failure to properly account for the many zeroes. Zika is generally considered more transmisable due to Ae. aegypti transmitting Zika at a higher rate (Freitas et al., 2019). This further illustrates it is important to account for zero inflation and to model the zero inflation properly (e.g., account for correlations).
@@
Covariates Zika-dengue chik.-dengue
Intercept 1.14 (1.03, 1.26) 1.02 (.91, 1.13)
$\text{verde}_{i}$ 1.02 (.95, 1.09) .92 (.85, 1)
$\text{SDI}_{i}$ 1.07 (.99, 1.15) 1.01 (.92, 1.11)
$\text{popdens}_{i}$ 1.02 (.94, 1.11) 1.06 (.96, 1.16)
$\text{favela}_{i}$98 (.91, 1.05) .94 (.87, 1.03)
$\text{temp}_{it}$ 1.14 (1.09, 1.20) .85 (.80, .90)
Neighborhood dengue prevalence .70 (.67, .74) .70 (.67, .74)
Neighborhood Zika prevalence 1.59 (1.48, 1.70)
Neighborhood chik. prevalence 1.43 (1.36, 1.50)
Previous Zika cases .90 (.85, .96)
Previous chik. cases .98 (.95, 1.02)
Interestingly, unlike in Schmidt et al. (2022), we did not find strong evidence that any of the spatial covariates were important for explaining differences in the transmission of the diseases. We found that Zika transmission was favored at warmer temperatures and chikungunya transmission was favored at colder temperatures, compared to dengue transmission. Both of these findings have been observed in laboratory studies (Tesla et al., 2018; Mercier et al., 2022). Indeed, a concern of chikungunya is that it could spread into colder areas of Europe unreachable by dengue (Mercier et al., 2022). To better quantify the temperature effects, we plot $\lambda_{kit}$, from Equation (2), versus temperature for $k=2$ (Zika) and $k=3$ (chikungunya) in Figure 2. From the figure, we have strong evidence, under otherwise average conditions, that Zika transmission was favored above 33.2 degrees and chikungunya transmission was favored below 32 degrees, compared to dengue transmission.
@@
The neighborhood prevalence effects are by far the strongest in Table 2 for explaining differences in the relative distribution of the diseases. For instance, we estimated that a standard deviation increase in the log of the prevalence of Zika across neighboring areas during the previous week was associated with a 59 (48, 70) percent increase in the odds of a new case being Zika relative to dengue. As Aedes aegypti primarily bite during the day, its arboviruses can spread rapidly between areas (Stoddard et al., 2013). We found that areas with more Zika cases were less favorable towards the transmission of chikungunya compared to dengue. That is, we estimated that a standard deviation increase in the log of the previous number of Zika cases in an area was associated with a 10 (4, 15) percent decrease in the odds of a new case being chikungunya relative to dengue. This does not necessarily mean that Zika transmission inhibited chikungunya transmission in a non-relative sense. It could be that Zika transmission encouraged both dengue and chikungunya transmission but the effect on dengue transmission was larger. Therefore, this result needs to be interpreted with some caution.
@@
We estimated $\zeta_{1}$, $\zeta_{2}$ and $\zeta_{3}$, from Equation (2), as .44 (.38, .49), .36 (.31, .32) and .43 (.36, .51) respectively, indicating a high degree of nonhomogeneous mixing. That is, the model predicts that if one of the diseases becomes very dominant in a neighborhood the relative distribution of the diseases will move towards a uniform distribution, perhaps because individuals avoid infection from the dominant disease. However, as avoiding one of the diseases should help avoid the others, as they are transmitted by the same vector, $\zeta_{k}$ could also be mainly reflecting differences in disease immunity, which are hard to account for explicitly, see SM Section 8. From Equations (3)-(4), we estimated the standard deviation of $\phi_{2it}$ as .75 (.71, .79), of $\phi_{3it}$ as .80 (.75, .86) and the correlation between $\phi_{2it}$ and $\phi_{3it}$ as .59 (.52, .65). This implies that the correlation between the disease cases, marginalizing the random effects and conditional on their total, is about what we would expect without the random effects, see the simulation study in SM Section 2. That is, the random effects do not induce any excess correlation between the disease counts. However, they do induce a large amount of overdispersion, see Figure 4.
@@
Table 3: Posterior means and 95% posterior credible intervals (in parentheses) for the estimated parameters from the Markov chain part of the fitted MS-ZIARMN model. The intercept row shows the probability of Zika presence and the logit of the probability of chikungunya presence, assuming no previous disease presence in the area, no cases in neighboring areas and at average values of the other covariates. The Zika column shows the ratio of $p_{2it}/(1-p_{2it})$ (odds of Zika presence) corresponding to a unit increase in the covariate. The chik. column shows the untransformed estimates (not exponentiated, see text). All covariates are standardized except the neighborhood prevalences which are given in units of 1 case per 25,000 (average combined size of neighboring areas) at small prevalences.
@@
Covariate Zika presence chik. presence
Intercept .1 (.08, .12) -9.06 (-11.97, -7)
$\text{verde}_{i}$ 1.14 (.98, 1.31) 08 (-.20 .37)
$\text{SDI}_{i}$ 1.11 (.96, 1.29) .12 (-.18, .42)
$\text{popdens}_{i}$ 1 (.85, 1.19) .02 (-.33, .37)
$\text{favela}_{i}$ 1.03 (.89, 1.19) 0 (-.3, .29)
$\text{temp}_{it}$ 1.44 (1.24, 1.66) 1.11 (.6, 1.7)
Neighborhood Zika prevalence 1.70 (1.55, 1.87)
Neighborhood chik. prevalence 1.09 (.85, 1.35)
Previous Zika presence 6.41 (3.95, 9.94) 5.12 (3.2, 8)
Previous chik. presence 3.46 (2.42, 4.98) 11.55 (8.13, 16.12)
Table 3 shows the estimated parameters from the Markov chain component of the model, that is, from Equation (7). The Zika column shows the ratio of the odds of Zika presence corresponding to a unit increase in the covariate. However, the intercept for chikungunya is highly negative making the odds ratios difficult to interpret and, therefore, we did not transform the chikungunya estimates. From the intercept row, chikungunya was very unlikely to emerge in an area under average conditions. Indeed, chikungunya emergence was only likely after Zika had emerged and temperatures rose in the next summer. In contrast, the emergence of Zika was much more random, occurring with a 10 (8, 12) percent chance bi-weekly under average conditions. Once chikungunya did emerge it was likely to stay in the area (probability of .88 (.63, .99) assuming no neighboring cases, no Zika presence in the previous bi-week and otherwise average conditions). Both Zika and chikungunya were more likely to be present at high temperatures and when there were cases in neighboring areas. The effects of the neighborhood prevalences are especially large, again illustrating the disease’s high rate of geographical spread. Zika presence was an indicator for chikungunya presence and chikungunya presence was an indicator for Zika presence. This could be unrelated to disease interactions and due to unmeasured confounding as the diseases share the same vector.
@@
The maps in Figure 3 show the posterior mean of ${\bar{\lambda}_{ki}=\mathopen{}\mathclose{{}\left(1/\sum_{t=2}^{T}S_{kit}}\right)\sum_{t=2}^{T}\lambda_{kit}S_{kit}}$ (average value of $\lambda_{kit}$ when disease $k$ was present in neighborhood $i$) for $k=2$ (Zika) (top map) and $k=3$ (chik.) (bottom map). From the bottom map, chikungunya transmission was disfavored in the west of the city when it was present, and the share of chikungunya cases tended to decline there. This was partly due to temperatures being 2-3 degrees higher on average in the west, see SM Figure 2, and chikungunya transmission being favored at colder temperatures relative to the other two diseases, see Table 2. Also, there were many Zika cases in the west of the city, and chikungunya transmission was disfavored in areas with many Zika cases. Zika transmission was favored throughout most of the city when it was present. This can be partly explained by the inherent favorability of Zika, i.e., the intercept in Table 2; temperatures typically being higher in Rio when Zika was present (the temperature covariate was centered using the overall average temperature) and Zika transmission being favored at higher temperatures.
@@
Figure 4: Top graphs show the posterior means (solid lines) and 95% posterior credible intervals (shaded areas) of the fitted values $y_{kit}^{\prime}$, see Section SM Section 5. The bottom graphs show the posterior probability the disease was present in each bi-week, $P(S_{kit}=1|\bm{y})$ versus $t$. "MS-ZIARMN no random effect" refers to the MS-ZIARMN model fit without including $\phi_{kit}$ from Equations (3)-(4).

### v2

@@ 4.1 Model specification and fitting
We modeled the Rio data with a bi-weekly time step as this roughly corresponds to the serial interval of the three diseases (Majumder et al., 2016; Riou et al., 2017). We took dengue as the baseline disease (k=1). Recall, from Section 2.1, that this means we are assuming that dengue was present in every neighborhood of Rio throughout the study period. We believe this is a reasonable assumption, as dengue has been circulating in Rio since 1986 and is well established there (Teixeira et al., 2009). We then considered $k=2$ to represent Zika and $k=3$ to represent chikungunya.
@@
For $\bm{x}_{kit}$ in (3), covariates that may affect the favorability of Zika or chikungunya transmission relative to dengue transmission, we first took all spatial covariates considered in Schmidt et al. (2022): the percentage of neighborhood $i$ covered in green area, $\text{verde}_{i}$; the social development index of neighborhood $i$, $\text{SDI}_{i}$; and the population density of neighborhood $i$, $\text{popdens}_{i}$. We additionally considered the percentage of neighborhood $i$ that was occupied by favelas (slums), $\text{favela}_{i}$. These spatial covariates were chosen as the transmission of arboviral diseases can be significantly influenced by socioeconomic factors such as water supply and urbanicity (Schmidt et al., 2011), see Schmidt et al. (2022) for more details. As for space-time covariates, we first took the average weekly maximum temperature in neighborhood $i$ during bi-week $t$, $\text{temp}_{it}$. Temperature is an important factor to consider, since if the transmission of Zika or chikungunya is more sensitive (less sensitive) to temperature compared to dengue, their range of spread will be smaller (bigger) (Tesla et al., 2018; Mercier et al., 2022). Note that if there are many more cases of Zika or chikungunya in neighboring areas compared to dengue, then we would expect the share of those diseases to grow, all else being equal, due to between-area mixing (Stoddard et al., 2013). Therefore, we considered the previous prevalence of disease $k$ and disease $1$ (dengue) across neighboring areas, $\log\mathopen{}\mathclose{{\left(\sum_{j\in NE(i)}y_{kj(t-1)}/\sum_{j\in NE(i)}\text{pop}_{j}+1}}\right)$ and $\log\mathopen{}\mathclose{{\left(\sum_{j\in NE(i)}y_{1j(t-1)}/\sum_{j\in NE(i)}\text{pop}_{j}+1}}\right)$, in $\bm{x}_{kit}$, where $NE(i)$ is the set of neighboring areas of neighborhood $i$ and $\text{pop}_{j}$ is the population of neighborhood $j$. We considered two areas neighbors if they shared a border. To simplify the model, we assumed the effect of neighboring dengue prevalence was the same on both $\log(\pi_{2it}/\pi_{1it})$ and $\log(\pi_{3it}/\pi_{1it})$, like with the within-area autoregression. Finally, it was speculated in Freitas et al. (2019) that Zika circulation could have been inhibiting chikungunya transmission. To account for these kinds of disease interactions, we considered $\log(y_{3i(t-1)}+1)$ in $\bm{x}_{2it}$ and $\log(y_{2i(t-1)}+1)$ in $\bm{x}_{3it}$. That is, chikungunya cases were allowed to affect the favorability of Zika transmission relative to dengue transmission, and Zika cases were allowed to affect the favorability of chikungunya transmission relative to dengue transmission.
@@
As for covariates potentially associated with the presence of Zika or chikungunya, $\bm{z}_{kit}$ in (7), we considered mostly the same factors included in $\bm{x}_{kit}$. An exception is that we did not include cases of the other diseases, as the Markov chain defined in (7) already accounts for disease interactions through $\rho_{jk}^{DI}$. Also, since the presence is not defined relatively, we did not include neighboring dengue prevalence.
@@
We also need to specify the initial state distributions, see Section 3. As there were no chikungunya cases reported at all in the first year, we assumed chikungunya had only a 5 percent chance of being present at the start of 2015. For Zika, there were a few, but not many, cases reported through most of 2015; therefore, we assumed a 10 percent chance of initial presence. We fitted the MS-ZIARMN model specified above to the Rio data using our proposed Gibbs sampler from Section 3. We ran the Gibbs sampler for 250,000 iterations on three chains with an initial burn-in of 50,0000 iterations. All sampling was started from random values in the parameter space to avoid convergence to local modes. Convergence was checked using the Gelman-Rubin statistic (<<1.05), the minimum effective sample size (>>1000) and by visually examining the traceplots (Plummer et al., 2006). The traceplots for all parameters (except the random effects) are shown in SM Section 10. It took around a day to run three chains of the Gibbs sampler in parallel on a laptop with an i9-12900HK, 2500 MHz, CPU.
@@
For comparison purposes, we also considered a model without zero inflation to ensure that there is justification for treating the many zeroes in the data, as hypothesized. That is, the ARMN model described in Equations (1)-(4) with the same covariates $\bm{x}_{2it}$ and $\bm{x}_{3it}$ as the MS-ZIARMN model. We also considered a zero-inflated model without the coupled Markov chains, that is, with $\rho_{k}^{AR}=0$ and $\rho_{jk}^{DI}=0$ for all $k$ and $j>1,\,j\neq k$ in Equation (7), which we will call the ZIARMN model. Without the coupled Markov chains, our model is a finite mixture and not a Markov mixture model, and the inference is greatly simplified, e.g., there is no need to use the FFBS algorithm. Finally, we compared with the zero-inflated multinomial model of Zeng et al. (2022) as our approach is built on their framework. Zeng et al. (2022) assumed that the probability a category was absent from an observation only depended on the category, so we let $\text{logit}(p_{kit})=\eta_{0k}$ for $k=2,3$ in Equation (7). We kept the rest of the model the same as the MS-ZIARMN model. Note that Zeng et al. (2022) did not consider any covariates or space-time structure in the multinomial probabilities. However, this is inappropriate for our example and did not fit the data well (results not shown). Also, Zeng et al. (2022) did not assume the baseline category was always present; however, this appears to be necessary when modeling a small number of categories, as discussed in Section 2.1.
@@
Table 1 shows the widely applicable information criterion (WAIC) (Gelman et al., 2014) of the 4 considered models. We calculated the WAIC by marginalizing out the state indicators $\bm{S}^{*}$ as recommended by Auger-Méthé et al. (2021), see SM Section 5 for more details. Note that the model with the smallest WAIC is considered to have the best fit, and as a rule of thumb, a difference of 5 or more in the WAIC is considered significant. Table 1 shows it is not only important to account for zero inflation, but also covariates and correlations across space, time and disease in the zero-inflated process (i.e., the process generating disease presence). Indeed, the model of Zeng et al. (2022) had a worse fit compared to the ARMN model (which does not account for zero-inflation), despite the other zero-inflated models having a much better fit.
@@ 4.2 Results
Table 2 shows the estimated coefficients from the multinomial part of the fitted MS-ZIARMN model, that is, the intercepts and covariate effects from Equation (3). From the intercept row, at average values of the covariates, Zika transmission was favored over dengue and chikungunya transmission. That is, under average conditions, the share of Zika cases tended to grow over time in an area (at least from an equal share, see the nuances in Section 2). It is interesting to compare these estimated intercepts with those from the ARMN and Zeng (2022) models in SM Table 1. Unlike the MS-ZIARMN model, the Zeng (2022) and ARMN models both estimated that, on average, Zika and chikungunya transmission was much less intense in Rio compared to dengue transmission. This does not correspond to our knowledge of the epidemiology of these diseases and is likely due to a failure to properly account for the many zeroes. Zika is generally considered more transmissible due to Aedes aegypti transmitting Zika at a higher rate (Freitas et al., 2019). This further illustrates it is important to account for zero inflation and to model the zero inflation properly (e.g., account for correlations).
@@
Covariates Parameter Zika-dengue chik.-dengue
Intercept $\exp({\alpha_{0}})$ 1.14 (1.03, 1.26) 1.02 (.91, 1.13)
$\text{verde}_{i}$ $\exp({\alpha_{\text{verde}}})$ 1.02 (.95, 1.09) .92 (.85, 1)
$\text{SDI}_{i}$ $\exp({\alpha_{\text{SDI}}})$ 1.07 (.99, 1.15) 1.01 (.92, 1.11)
$\text{popdens}_{i}$ $\exp({\alpha_{\text{popdens}}})$ 1.02 (.94, 1.11) 1.06 (.96, 1.16)
$\text{favela}_{i}$ $\exp({\alpha_{\text{favela}}})$ .98 (.91, 1.05) .94 (.87, 1.03)
$\text{temp}_{it}$ $\exp({\alpha_{\text{temp}}})$ 1.14 (1.09, 1.20) .85 (.80, .90)
Neighborhood dengue prevalence $\exp({\alpha_{\text{NeiDeng}}})$ .70 (.67, .74) .70 (.67, .74)
Neighborhood Zika prevalence $\exp({\alpha_{\text{NeiZika}}})$ 1.59 (1.48, 1.70)
Neighborhood chik. prevalence $\exp({\alpha_{\text{NeiChik}}})$ 1.43 (1.36, 1.50)
Previous Zika cases $\exp({\alpha_{\text{Zika}}})$ .90 (.85, .96)
Previous chik. cases $\exp({\alpha_{\text{Chik}}})$ .98 (.95, 1.02)
Interestingly, unlike in Schmidt et al. (2022), we did not find strong evidence that any of the spatial covariates were important for explaining differences in the transmission of the diseases. We found that Zika transmission was favored at warmer temperatures and chikungunya transmission was favored at colder temperatures, compared to dengue transmission. Both of these findings have been observed in laboratory studies (Tesla et al., 2018; Mercier et al., 2022). Indeed, a concern with chikungunya is that it could spread into colder areas of Europe that are unreachable by dengue (Mercier et al., 2022). To better quantify the temperature effects, we plot $\lambda_{kit}$, from Equation (2), versus temperature for $k=2$ (Zika) and $k=3$ (chikungunya) in Figure 2. From the figure, we have strong evidence that, under otherwise average conditions, Zika transmission was favored above 33.2 degrees and chikungunya transmission was favored below 32 degrees, compared to dengue transmission (the average temperature in Rio during the study was 33.8 degrees, with the middle 95 percent being from 27.2 to 39.1 degrees).
@@
The neighborhood prevalence effects are by far the strongest in Table 2 for explaining differences in the relative distributions of the diseases. For instance, we estimated that a standard deviation increase in the log of the prevalence of Zika across neighboring areas during the previous week was associated with a 59 (48, 70) percent increase in the odds of a new case being Zika relative to dengue. As Aedes aegypti primarily bite during the day, its arboviruses can spread rapidly between areas (Stoddard et al., 2013). We found that areas with more Zika cases were less favorable towards the transmission of chikungunya relative to dengue. That is, we estimated that a standard deviation increase in the log of the previous number of Zika cases in an area was associated with a 10 (4, 15) percent decrease in the odds of a new case being chikungunya relative to dengue. This does not necessarily mean that Zika transmission inhibited chikungunya transmission in a non-relative sense. It could be that Zika transmission encouraged both dengue and chikungunya transmission, but the effect on dengue transmission was larger. Therefore, this result needs to be interpreted with some caution.
@@
We estimated $\zeta_{1}$, $\zeta_{2}$ and $\zeta_{3}$ from Equation (2) as .44 (.38, .49), .36 (.31, .32) and .43 (.36, .51), respectively, indicating a high degree of nonhomogeneous mixing. That is, the model predicts that if one of the diseases becomes very dominant in a neighborhood, the relative distribution of the diseases will move towards a uniform distribution, perhaps because individuals avoid infection from the dominant disease. However, as avoiding one of the diseases should help avoid the others, as they are transmitted by the same vector, $\zeta_{k}$ could also be mainly reflecting differences in disease immunity, which are difficult to account for explicitly (see SM Section 9). From Equations (3)-(4), we estimated the standard deviation of $\phi_{2it}$ as .75 (.71, .79), of $\phi_{3it}$ as .80 (.75, .86) and the correlation between $\phi_{2it}$ and $\phi_{3it}$ as .59 (.52, .65). This implies that the correlations between the disease cases, marginalizing the random effects and conditional on their total, is about what we would expect without the random effects, see the simulation study in SM Section 2. That is, the random effects do not induce any excess correlations between the disease counts. However, they do induce a large amount of overdispersion, as seen in Figure 4.
@@
Table 3: Posterior means and 95% posterior credible intervals (in parentheses) for the estimated parameters from the Markov chain part of the fitted MS-ZIARMN model. The intercept row shows the probability of Zika presence and the logit of the probability of chikungunya presence, assuming no previous disease presence in the area, no cases in neighboring areas and at average values of the other covariates. The Zika column shows the ratio of $p_{2it}/(1-p_{2it})$ (odds of Zika presence) corresponding to a unit increase in the covariate. The chik. column shows the untransformed estimates (not exponentiated, see text). All covariates are standardized except the neighborhood prevalences, which are given in units of 1 case per 25,000 (average combined size of neighboring areas) at small prevalences.
@@
Covariate Parameter Zika presence chik. presence
$\text{Zika col.}\,\,\,|\,\,\,\text{chik. col.}$
Intercept $\exp(\eta_{0})/(1+\exp(\eta_{0}))\,\,\,|\,\,\,\eta_{0}$ .1 (.08, .12) -9.06 (-11.97, -7)
$\text{verde}_{i}$ $\exp(\eta_{\text{verde}})\,\,\,|\,\,\,\eta_{\text{verde}}$ 1.14 (.98, 1.31) 08 (-.20 .37)
$\text{SDI}_{i}$ $\exp(\eta_{\text{SDI}})\,\,\,|\,\,\,\eta_{\text{SDI}}$ 1.11 (.96, 1.29) .12 (-.18, .42)
$\text{popdens}_{i}$ $\exp(\eta_{\text{popdens}})\,\,\,|\,\,\,\eta_{\text{popdens}}$ 1 (.85, 1.19) .02 (-.33, .37)
$\text{favela}_{i}$ $\exp(\eta_{\text{favela}})\,\,\,|\,\,\,\eta_{\text{favela}}$ 1.03 (.89, 1.19) 0 (-.3, .29)
$\text{temp}_{it}$ $\exp(\eta_{\text{temp}})\,\,\,|\,\,\,\eta_{\text{temp}}$ 1.44 (1.24, 1.66) 1.11 (.6, 1.7)
Neighborhood Zika prevalence $\exp(\eta_{\text{NeiZika}})\,\,\,|\,\,\,\eta_{\text{NeiZika}}$ 1.70 (1.55, 1.87)
Neighborhood chik. prevalence $\exp(\eta_{\text{NeiChik}})\,\,\,|\,\,\,\eta_{\text{NeiChik}}$ 1.09 (.85, 1.35)
Previous Zika presence $\exp(\rho_{2}^{AR})\,\,\,|\,\,\,\rho_{23}^{DI}$ 6.41 (3.95, 9.94) 5.12 (3.2, 8)
Previous chik. presence $\exp(\rho_{32}^{DI})\,\,\,|\,\,\,\rho_{3}^{AR}$ 3.46 (2.42, 4.98) 11.55 (8.13, 16.12)
Table 3 shows the estimated parameters from the Markov chain component of the model, that is, from Equation (7). The Zika column shows the ratio of the odds of Zika presence corresponding to a unit increase in the covariate. However, the intercept for chikungunya is highly negative, making the odds ratios difficult to interpret. Therefore, we did not transform the chikungunya estimates. From the intercept row, it was very unlikely for chikungunya to emerge in an area under average conditions. Indeed, chikungunya emergence was only likely after Zika had emerged and temperatures rose in the next summer. In contrast, the emergence of Zika was much more random, occurring with a 10 (8, 12) percent chance bi-weekly under average conditions. Once chikungunya did emerge, it was likely to stay in the area (probability of .88 (.63, .99) assuming no neighboring cases, no Zika presence in the previous bi-week and otherwise average conditions). In contrast, Zika persistence was more reliant on the presence of cases in neighboring areas (there was only a 41 (32, 49) percent chance of Zika being present if it was present previously and there were no neighboring cases). Both Zika and chikungunya were more likely to be present at high temperatures and when there were cases in neighboring areas. The effects of neighborhood prevalence are especially large, again illustrating the diseases’ high rate of geographical spread. The presence of Zika was an indicator of chikungunya presence, and vice versa. This could be unrelated to disease interactions and due to unmeasured confounding, as the diseases share the same vector.
Some of the chikungunya estimates in Table 3 are extreme. Chikungunya cases were first reported in most neighborhoods between bi-weeks 28 and 34 (measured from the beginning of 2015). This implies a significant change in the probability of chikungunya presence around bi-week 28 across all neighborhoods. Therefore, the extreme estimates in Table 3 are likely due to the model attempting to fit this large and fast shift in the probability of chikungunya presence. In light of the above, the chikungunya estimates in Table 3 should be interpreted with caution, as there may be unmeasured confounders driving the change in the risk of chikungunya presence around bi-week 28. For example, spread from travel outside the city (Nunes et al., 2015), which is difficult to measure and may be related to factors such as temperature. In contrast, Zika cases emerged much more randomly across the city, as reflected in Table 3.
@@
The maps in Figure 3 show the posterior mean of $\bar{\lambda}_{ki}=\mathopen{}\mathclose{{\left(1/\sum_{t=2}^{T}S_{kit}}}\right)\sum_{t=2}^{T}\lambda_{kit}S_{kit}$} (average value of $\lambda_{kit}$ when disease $k$ was present in neighborhood $i$) for $k=2$ (Zika) (top map) and $k=3$ (chik.) (bottom map). From the bottom map, chikungunya transmission was disfavored in the west of the city when it was present, and the share of chikungunya cases tended to decline there. This was partly due to temperatures being 2-3 degrees higher on average in the west (see SM Figure 2) and to chikungunya transmission being disfavored at warmer temperatures relative to the other two diseases (see Table 2). Also, there were many Zika cases in the west of the city, and chikungunya transmission was disfavored in areas with many Zika cases. Zika transmission was favored throughout most of the city when it was present. This can be partly explained by the inherent favorability of Zika, i.e., the intercept in Table 2; by temperatures typically being higher in Rio when Zika was present (the temperature covariate was centered using the overall average temperature); and by Zika transmission being favored at higher temperatures.
@@
Figure 4: Top graphs show the posterior means (solid lines) and 95% posterior credible intervals (shaded areas) of the fitted values $y_{kit}^{\prime}$, see SM Section 6. The bottom graphs show the posterior probability that the disease was present in each bi-week, $P(S_{kit}=1|\bm{y})$ versus $t$. "MS-ZIARMN no random effect" refers to the MS-ZIARMN model fit without including $\phi_{kit}$ from Equations (3)-(4).